OpenStudy (anonymous):

n 1977, the price of a scientific calculator was $175. In 2004, the price was $15. Assuming the change in price was linear, what was the approximate price of a scientific calculator in 1998? $23.00 $27.00 $51.00 $60.00

OpenStudy (anonymous):

repeat this one

OpenStudy (amistre64):

2004,15 1977,175 ---------- 27, -160; slope = y/x = -160/27 y = -(160/27)(x-2004)+15

OpenStudy (amistre64):

51

OpenStudy (anonymous):

thanks love (:

OpenStudy (amistre64):

:) my quad finder doesnt do well with lines lol. Youre welcome ;)